cos(15°) can be evaluated using a sum or difference trig identity. Which of the following is equivalent to cos(15°)?
1
Expert's answer
2020-11-16T14:03:06-0500

Let's apply the cos of a sum formula to "\cos(15^\circ+15^\circ)", as we know that "\cos(30^{\circ})=\frac{\sqrt{3}}{2}:"

"\frac{\sqrt{3}}{2} = \cos(30^\circ)=\cos(15^\circ+15^\circ)=\cos(15^\circ)\times\cos(15^\circ)-\sin(15^\circ)\times\sin(15^\circ)"

Now by the main trigonometrical identity ("\sin^2\alpha+\cos^2\alpha=1)" we can write :

"\frac{\sqrt{3}}{2}=\cos^2(15^\circ)-(1-\cos^2(15^\circ))=2\cos^2(15^\circ)-1"

"2\cos^2(15^\circ)=1+\frac{\sqrt{3}}{2}"

"\cos(15^\circ)=\pm \sqrt{\frac{1}{2}+\frac{\sqrt{3}}{4}}"

We know also that "\cos(15^\circ)>0" (as "0^\circ<15^\circ<90^\circ" ), so we have :

"\cos(15^\circ)=\sqrt{\frac{1}{2}+\frac{\sqrt{3}}{4}}=\frac{1}{2} \sqrt{2+\sqrt{3}}"

We can also simplify the last expression by writing :

"(x+y)^2=2+\sqrt{3}"

"x^2+y^2=2, xy=\frac{\sqrt{3}}{2}"

"x^2+\frac{3}{4x^2}=2"

"x^4-2x^2+\frac{3}{4}=0"

"x^2=\frac{2\pm\sqrt{4-3}}{2}"

"x=\sqrt{\frac{1}{2}},y=\sqrt{\frac{3}{2}}"

So we have :

"\cos(15^\circ)=\frac{1}{2}\times(\frac{1}{\sqrt{2}}+\frac{\sqrt{3}}{\sqrt{2}})=\frac{\sqrt{2}+\sqrt{6}}{4}"